OpenStudy (anonymous):

simplify 8 divided by (1 1/4 divided by 5)

OpenStudy (anonymous):

Is this written as \[8\div\frac{ 1 \frac{ 1 }{ 4 } }{5 }\] ?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

you have to do the parenthesis first

OpenStudy (anonymous):

Dividing by a fraction is the same as multiplying by its reciprocal. Therefore, all you have to do is \[8 \times \frac{ 5 }{ 1\frac{ 1 }{ 4 } }\]

OpenStudy (anonymous):

Which is just \[ \frac{ 40 }{ 1 \frac{ 1 }{ 4 }}\]

OpenStudy (anonymous):

or \[\frac{ 40 }{ 1.25 }\]

OpenStudy (anonymous):

Or 32... Basic math.

OpenStudy (anonymous):

Thank you for your help